The following is the sales budget for Shleifer, Inc., for the first quarter of 2019:

January

February

March

  Sales budget

$

203,000

$

223,000

$

246,000

Credit sales are collected as follows:

65 percent in the month of the sale.

20 percent in the month after the sale.

15 percent in the second month after the sale.

The accounts receivable balance at the end of the previous quarter was $87,000 ($57,000 of which were uncollected December sales).

a.         Calculate the sales for November. (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)

b.        Calculate the sales for December. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

c.         Calculate the cash collections from sales for each month from January through March. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
Answer #1
a.
Uncollected November sales at the end of previous quarter = Accounts receivables balance at the end of previous quarter - Uncollected December sales = 87000 - 57000   30000
Sales for November = Uncollected November sales at the end of previous quarter / % Uncollected = 30000 / 15% 200000
b.
Sales for December = Uncollected December sales at the end of previous quarter / % Uncollected = 57000 / 35% 162857.14
c.
Cash collections from sales : January February March

65% of current month sales

131950

[ 65%*203000 ]

144950

[ 65%*223000 ]

159900

[ 65%*246000 ]

20% of previous months sales

32571.43

[ 20%*162857.14 ]

40600

[ 20%*203000 ]

44600

[ 20%*223000 ]

15% of sales made 2 month ago

30000

[ 15%*200000 ]

24428.57

[ 15%*162857.14 ]

30450

[ 15%*203000 ]

Cash collections 194521.43 209978.57 234950.00